OpenStudy (anonymous):

Which expression is equivalent to 6a + 15? A. 6(a + 5) B. 3(2a + 5) C. 3(3a + 12) D. 6(a + 12) PLEASE HELP :)

jimthompson5910 (jim_thompson5910):

with choice A, we can distribute to get 6(a + 5) = 6*a + 6*5 = 6a + 30 the result we get above is NOT the same as 6a+15, so choice A is out

jimthompson5910 (jim_thompson5910):

do the same for the other choices

OpenStudy (anonymous):

how do i do it ?

jimthompson5910 (jim_thompson5910):

notice how I multiplied the outer 6 by each term inside the parenthesis 6 times a = 6a 6 times 5 = 30 so that's why 6(a+5) = 6a+30 making sense?

OpenStudy (anonymous):

6a+15=3*2a+3*5=3(2a+5)